OpenStudy (anonymous):

solve the following equation -4(2x-6)= 2(15-5x)

OpenStudy (anonymous):

Distribute first... post the answer to the distribute part and I will help you from there.

OpenStudy (anonymous):

A. -1/3 B. -2 C. -36/13 D. -6/13

OpenStudy (anonymous):

-8x + 24 = 30 - 10x add 10x to both sides. 2x + 24 = 30 subtract 24 2x = 6 divide by 2 x = 3

OpenStudy (anonymous):

so whats the answer? lol

OpenStudy (anonymous):

The answer is 3 all of your answers are incorrect if you copied them down correctly. If the problem says to solve. If it says something else then you need to post it.

OpenStudy (anonymous):

everything i wrote is correct as far as this paper has it down.

OpenStudy (anonymous):

Does it say to solve?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Well... 3 is the answer... You will just have to tell your teacher that none of the answers are correct.